If a,b,c are the sides of a Δ ABC such that x2-2(a+b+c)x +3Κ(ab+bc+ac)=0 has real roots then,


a) K <4/3


b) K>5/3


c) K ε ( 4/3,5/3)

Grade:12

1 Answers

saket shrivastava
12 years ago

the ans is  k<2/3

let a=b=c

option a==> let k=1 (<4/3),, the d=B2-4AC<0 not real

option b ===> let k=2 (>5/3) again d<0

option c ====> let k= 3/2  (={4/3+5/3}/2) again d<0

 

check options pls and approve